The probability that Ato , Sulley and Kofi will gain admission
to a certain university are 4/5,3/4 and 2/3 respectively.
Find the probability that :
(a) none of them will gain admission
(b) at most two of them will gain admission
(c) only Kofi and Ato will gain admission.

Answer by Edwin McCravy(20064)

The probability that Ato , Sulley and Kofi will gain admission
to a certain university are 4/5,3/4 and 2/3 respectively.
P(A) = 4/5, P(not A) = 1 - 4/5 = 1/5
P(S) = 3/4, P(not S) = 1 - 3/4 = 1/4
P(K) = 2/3, P(not K) = 1 - 2/3 = 1/3

Find the probability that :
(a) none of them will gain admission
P(not A)×P(not S)×P(not K) = (1/5)(1/4)(1/3) = 1/60

(b) at most two of them will gain admission
P(not all three) = 1 - P(all three) = 1 - P(A)×P(S)×P(K) =
1 - (4/5)(3/4)(2/3) = 1 - 2/5 = 3/5

(c) only Kofi and Ato will gain admission.
P(K)×P(A)×P(not S) = (2/3)(4/5)(1/4) = 2/15

Edwin